Consider a parallelogram ABCD. Show that AC2 + BD2 = AB2 + BC2 + CD2 + AD2


1
Expert's answer
2021-09-28T00:07:50-0400

Let us consider a parallelogram "ABCD" and show that "AC^2 + BD^2 = AB^2 + BC^2 + CD^2 + AD^2".

According to Cosine Theorem, "AC^2=AB^2+BC^2-2AB\cdot BC\cos\angle ABC" and "BD^2=BC^2+CD^2-2BC\cdot CD\cos\angle BCD."

Taking into acount then in the parallelogram "ABCD" the angles "\angle ABC" and "\angle BCD" are supplementary, and "AB=CD,\ BC=AD," we conclude that

"AC^2+BD^2=(AB^2+BC^2-2AB\cdot BC\cos\angle ABC)+(BC^2+CD^2-2BC\cdot CD\cos\angle BCD)\\
=AB^2+BC^2-2AB\cdot BC\cos\angle ABC+AD^2+CD^2-2BC\cdot AB\cos(180^{\circ}-\angle ABC)\\
=AB^2+BC^2+CD^2+AD^2-2AB\cdot BC\cos\angle ABC+2BC\cdot AB\cos(\angle ABC)\\
=AB^2+BC^2+CD^2+AD^2."